# Code Difference Guided Fuzzing for FPGA Logic Synthesis Compilers via Bayesian Optimization

### v1

@@ III-C Bayesian Diversity Selection Component
The Bayesian diversity selection component (as shown as Algorithm2) first convert the CPS model variant generated by the Test-Program Generation component into HDL Code through HDL Coder. Secondly, the Bayesian diversity selection component selects program variants with higher diversity through program distance and Bayesian selector. Through the Bayesian diversity selection component LSC-Fuzz, more diverse program variants are selected for testing, which enriches the diversity of the program and increases the ability of test cases to trigger deeper bugs.
@@
Input: Seed CPS model $c$, HDL Coder $L$, List of program variants $V$, Prior probabilities
@@ 10    $d(V_{i})\leftarrow\text{FDist}(H,V_{i})$;
The Bayesian selector is a decision-making tool based on Bayesian decision theory and is used to make optimal choices under uncertainty. It uses prior knowledge and newly acquired data to update probability estimates to guide decision-making. The basic strategy of Bayesian selectors is to select variants with high temporal complexity and program complexity. LSC-Fuzz first define the program distance $d$ between two programs as follow.
@@
Then LSC-Fuzz utilize Bayesian inference. Prior probability $P(V_{i})$ for each variant $V_{i}$ represents our initial belief about its complexity. The likelihood $P(D|V_{i})$, indicates the probability of observing specific complexity metrics $D$ given a variant $V_{i}$. The posterior probability $P(V_{i}|D)$, which updates our belief in the complexity of each variant based on observed data, is computed using formula (2):
@@
In this way, LSC-Fuzz calculate the acceptance probability $P(V_{i}|D)$ of all variant $V_{i}$ and form these probabilities into a one-dimensional matrix. This simplifies the selection of variants. If the test program setting requires multiple variants, LSC-Fuzz select those with the highest acceptance probability from the matrix. Through Bayesian selectors, LSC-Fuzz efficiently generates more complex HDL code for variant test programs.
@@ III-D Equivalent Check Component
Since the netlist files generated by logic synthesis are usually not directly comparable, LSC-Fuzz have chosen two verification methods to verify the consistency of the netlist file, Satisfiability Modulo Theories(SMT) formalize verification and simulation verification.
III-D1 SMT formalize verification
Formal verification with SMT solvers entails several steps, ranging from defining logical constraints to solving them. LSC-Fuzz compared different netlists using the SMT solver ABC, provided by Yosys. The first step in SMT verification is to clarify the constraints.
$\mathcal{P}(\mathcal{N})=\{\phi_{1},\phi_{2},\ldots,\phi_{n}\}$ (4)
Fomula (4) introduces $\mathcal{P}$, which represents the set of properties $\phi_{i}$ associated with a netlist $\mathcal{N}$. These properties include functional characteristics and constraints that the netlist must satisfy.
After defining the constraints, LSC-Fuzz use the solver to check whether the constraints are satisfied, as described in Formula (5).
$\mathcal{C}(\mathcal{N}_{\text{orig}},\mathcal{N}_{\text{var}})=\bigwedge_{i=1}^{n}(\phi_{i}^{\text{orig}}\leftrightarrow\phi_{i}^{\text{var}})$ (5)
where $\mathcal{C}$ denotes the conjunction of bi-conditional constraints between corresponding properties of the original $\mathcal{N}_{\text{orig}}$ and variant $\mathcal{N}_{\text{var}}$ netlists, ensuring that each property $\phi_{i}$ is equivalently satisfied in both.
The SMT solver performs these checks to confirm that the netlist file is consistent before and after simulation. If there is an inconsistency, LSC-Fuzz will report the problem.
Figure 6: The output of FPGA logic synthesis compilers
It is worth noting that although the variant file generated by LSC-Fuzz is theoretically consistent with the output of the original file. However, they are not completely consistent in the SMT equivalence check, because the parts that will not be executed will also be reflected in the netlist file. Therefore, the SMT check has a false positive problem when verifying the variant file and the original file. To solve this problem, LSC-Fuzz chose to use simulation verification to further verify the FPGA logic synthesis compiler
III-D2 Simulation verification
This component consists of two sub-processes using different simulation tools and comparing original and variant model outputs. The output of the simulation tool is a waveform containing high and low levels(as shown in the Figure 6), so LSC-Fuzz print the output as a 0/1 level signal format and compare them bit by bit.
For different simulation tools, LSC-Fuzz use the same file to generate netlist files through different FPGA logic synthesis compilers, and use the same stimulus file to simulate them to compare their outputs. Fomula (6) represents the execution of simulation $\mathcal{S}_{j}$ on netlist $\mathcal{N}$ using stimuli $\sigma$, resulting in output $\mathcal{O}_{j}$. Each simulation tool might use a slightly different engine or algorithm, and $\mathcal{O}_{j}$ captures the output of the $j$-th tool.
@@ III-D2 Simulation verification
After simulations, outputs are compared to assess the consistency of the netlist’s behavior across different simulation platforms based on Fomula (7). This formula ensures that all outputs $\mathcal{O}_{i}$ are identical for every pair of simulations, confirming that the netlist behaves consistently irrespective of the simulation tool used. If the output is inconsistent, LSC-Fuzz will report the issue. We will manually check the issues reported by LSC-Fuzz and report to developer.
@@
For comparing original and variant model outputs, LSC-Fuzz prints the signal as a 0/1 electrical signal, and compares the simulation result output after the original CPS model is converted to HDL code and synthesized bit by bit with the simulation result output after the variant CPS model is converted into HDL code and synthesized. Theoretically the results should not change since the inserted module will not be executed and translated. But if the results are inconsistent, LSC-Fuzz will report the vulnerability. We will manually check these issues to find the root causes and submit them to the corresponding software developers.
@@ IV Evaluation
In this section, four experiments are conducted to evaluate the effectiveness of LSC-Fuzz. Specifically, our evaluation aims at answering the following Research Questions (RQs).
RQ1: How is the bug-finding capability of LSC-Fuzz compared to the state-of-the-art methods?
RQ2: Can LSC-Fuzzdetect new FPGA logic synthesis compiler bugs?
RQ3: How effectiveness is the HDL code generation of LSC-Fuzz compared to the state-of-the-art methods?
RQ4: How effectiveness is the Bayes Selector of LSC-Fuzz?
In our experiments, RQ1 and RQ2 are used to evaluate the bug-finding capability of LSC-Fuzz compared to the state-of-the-art approaches. RQ3 and RQ4 are employed to evaluate the code generation strategy of the code generation component and the diverse variant generation component.
@@ IV-B Evaluation Setup
LSC-Fuzz has been developed using MATLAB and Python, and both our code and experimental data are publicly accessible on GitHub [9]. The evaluation of LSC-Fuzz was conducted on a computer running the Ubuntu 22.04 operating system, equipped with an Intel Core i9 CPU @ 2.10GHz, and 128GB of memory. Based on the recommendations of existing work, we focus on the size of generated HDL files being in the 700-100 line range [7].
@@ IV-C Answer to RQ1
Approach. To evaluate the effectiveness of LSC-Fuzz, we compare the bug-finding capability of LSC-Fuzz with the state-of-the-art methods Verismith [7] and LegoHDL [8], since finding more bugs within a time period is the main objective of these approaches. In the experiment, we detect bugs on both the recently released version of FPGA logic synthesis compiler include Vivado, yosys, Iverilog and Quartus. We set a single testing period of two weeks for FPGA logic synthesis compiler, that is, every method tests these FPGA sythesis compilers for two weeks.
TABLE I: Bugs detected by LSC-Fuzz, LegoHDL, and Verismith
@@
LSC-Fuzz 2 4 6
Results. As illustrated in Table I, bugs detected in our experiment are categorized into new bugs (New) and previously identified ones (Known), with the latter being duplicates from the bug repository. Notably, Table I demonstrates that LSC-Fuzz significantly outperforms Verismith and LegoHDL in bug detection capabilities. Over four weeks, LSC-Fuzz identified six bugs in these FPGA logic synthesis compilers, comprising two new bugs and four known ones. In contrast, LegoHDL detected three bugs—two known and one new—while Verismith identified only one known bug. The comparative analysis of bugs detected by Verismith, LegoHDL, and LSC-Fuzz in FPGA logic synthesis compilers is detailed in Figure 7. This occurs because LSC-Fuzz is capable of generating HDL test cases that are more complex than those produced by the baseline method. During the translation of the CPS model to HDL code by the HDL coder, each subsystem is converted into a separate HDL file with defined reference relationships. LSC-Fuzz enhances the complexity of these relationships by integrating complex code branches into the CPS model, thereby challenging the FPGA logic synthesis compiler to employ more sophisticated optimization strategies for the HDL files. Additionally, the Bayesian selector intensively tests the compiler by selecting increasingly complex HDL code scenarios for processing. All these lead to LSC-Fuzz’s bug discovery efficiency being better than the baseline method.
Conclusion. LSC-Fuzz significantly outperforms LegoHDL and Verismith for testing FPGA logic synthesis compiler. These results verify the bug-finding capability of LSC-Fuzz.
@@ IV-D Answer to RQ2
Approach. Over a two-month period from May to July 2024, we extensively tested LSC-Fuzz to assess its bug-finding capabilities. Unfortunately, we could not validate the latest version of Quartus due to licensing issues. However, we employed the latest versions of other FPGA logic synthesis compilers for testing. Each identified issue with a logic synthesis compiler was reported to the corresponding software community, and we engaged with developers there to gather feedback.
Results. As depicted in Table II, our approach successfully identified 16 bugs within a three-month timeframe. The official websites of various logic synthesis tools, including Iverilog, Vivado, Yosys, and Quartus, have classified each issue as either Confirmed as New, Confirmed as Known, or Unconfirmed in Doubt (Pending Verification). New indicates bugs previously unrecognized by developers. Known refers to issues already identified as bugs by developers. Pending represents issues that developers believe can be circumvented through specific standardized procedures. To facilitate replication of our findings, we have made the HDL code files that triggered these bugs available on GitHub [9]. Notably, 4 new bugs were detected, which were not identified by previous methods. Figure 8 showcases the bug discovered by LSC-Fuzz. During the synthesis of HDL code, the compiler failed to correctly manage the reference relationships, leading to incorrect execution within the module positive. The Nonpositive block, correctly transformed into Verilog, reveals how intricate reference relationships can induce synthesis errors and potential compiler crashes. Existing methods like Verismith do not detect this vulnerability due to their limited code generation scopes that fail to produce complex reference relationships between blocks. This bug significantly impacts Vivado [10].
TABLE II: Vulnerability found by LSC-Fuzz
@@
Vivado-2 Synthesis faild and crush(HARTGLADDGen:regenerate(bool)) Verified C Vivado
Vivado-3 Vivado synthesis faild (ConstPropagte) Verified C Vivado
@@
Vivado-7 Reference relationship error casued synthesis failed Pending C Vivado
@@
    There are two types of status feedback from MathWorks on bug report (i.e., $Verified$ = newly confirmed bug, pendingpending = Pending verification). There are two types of bugs (i.e., Type) in our reported bugs: crash bugs ($C$) and miscompilation bugs ($M$).
Conclusion. LSC-Fuzz has proven effective in detecting bugs in FPGA logic synthesis compilers. Within just two months, we reported 16 valid bugs, four of which have been newly confirmed.
@@ IV-E Answer to RQ3
Approach. We employ LSC-Fuzz, LegoHDL, and Verismith to generate HDL code test cases, each consisting of 700-1000 lines and totaling 1000 test cases. We compare their complexity and generation time using three key indicators derived from the Abstract Syntax Tree (AST) provided by Yosys. Our first metric, the count of blocks, involves tallying the blocks in files generated by both our model and the baselines, with a higher block count indicating greater complexity. The second metric, the count of lines or connections, suggests that a higher number of connections might increase the likelihood of encountering bugs. The third metric evaluates specific reference relationships, which when complex, are more likely to expose bugs.
Figure 7: Real logic synthesis compiler bug found by LSC-Fuzz, LegoHDL and Verismith
Figure 8: Real lofic synthesis compiler bug found by LSC-Fuzz
Results. Numbers of Blocks and Connections. Blocks and connections are vital components of HDL code, crucial for depicting its characteristics. As shown in Figure9, LSC-Fuzz generates a range of [40,49] blocks and [125,226] connections, surpassing the ranges produced by Verismith [7] ([25,37] blocks and [102,117] connections) and LegoHDL ([30,41] blocks and [120,210] connections). This increased complexity arises as LSC-Fuzz continuously selects more intricate HDL codes for testing, based on a Bayesian selector, enhancing its bug detection efficiency over the baseline methods. Specific Reference Relationships. In HDL design, complex reference relationships can challenge the compiler’s execution capabilities and optimize circuit design. To quantify this metric, we fixed the top-level design module and counted the number of modules it references. Figure9 shows that LSC-Fuzz’s range of module references exceeds those of Verismith [7] and LegoHDL. Specifically, LSC-Fuzz’s median reference count is 20% higher than LegoHDL’s and 33.3% higher than Verismith’s. On average, LSC-Fuzz’s reference count is 16.67% higher than LegoHDL’s and 25% higher than Verismith’s. The increase is attributed to LSC-Fuzz’s addition of more If branches, which are translated into independent files with reference relationships during translation, thereby complicating the application relationship. This complexity explains LSC-Fuzz’s ability to uncover issues that baseline methods miss, as it more thoroughly exercises the FPGA logic synthesis compiler with these complex relationships.
Conclusion. LSC-Fuzz outperforms baseline methods by generating more complex HDL test cases, enhancing its capability to detect bugs in FPGA logic synthesis compilers.

### v2

@@ III-C Bayesian Diversity Selection Component
The BDS first convert the Simulink Model variant generated by the Test-Program Generation component into HDL Code through HDL Coder.Secondly, we recognize that the inherent optimizations of HDL Coder may cause different Simulink Models to generate structurally similar HDL programs, which reduces the effectiveness of testing. To address this issue, the Bayesian Diversity Selection (BDS) component selects program variants with higher structural diversity by combining program distance with a Bayesian selector.
@@
Input: Seed Simulink Model $c$, HDL Coder $L$, List of program variants $V$, Prior probabilities
@@ 10    $d(V_{i})\leftarrow\text{FDist}(H,V_{i})$;
The Bayesian selector is a decision-making tool based on Bayesian decision theory and is used to make optimal choices under uncertainty. It uses prior knowledge and newly acquired data to update probability estimates to guide decision-making. The basic strategy of Bayesian selectors is to select variants with high temporal complexity and program complexity. SmootHDL first define the program distance $d$ between two programs as follow.
@@
Then SmootHDL utilize Bayesian inference. Prior probability $P(V_{i})$ for each variant $V_{i}$ represents our initial belief about its complexity. The likelihood $P(D|V_{i})$, indicates the probability of observing specific complexity metrics $D$ given a variant $V_{i}$. The posterior probability $P(V_{i}|D)$, which updates our belief in the complexity of each variant based on observed data, is computed using formula (9):
@@
In this way, SmootHDL calculate the acceptance probability $P(V_{i}|D)$ of all variant $V_{i}$ and form these probabilities into a one-dimensional matrix. This simplifies the selection of variants. If the test program setting requires multiple variants, SmootHDL select those with the highest acceptance probability from the matrix. Through Bayesian selectors, SmootHDL efficiently generates more complex HDL code for variant test programs.
@@ III-D Equivalent Check Component
Since the netlist files generated by logic synthesis are usually not directly comparable, SmootHDL has chosen simulation to verify the consistency of the netlist file. In addition, Yosys provides the abc engine for equivalence checking of netlist files, which we use as an auxiliary method to support verification, although it is not the primary approach in our framework.
The ETC consists of two sub-processes using different simulation tools and comparing original and variant model outputs. The output of the simulation tool is a waveform containing high and low levels(as shown in the Figure 3), so SmootHDL print the output as a 0/1 level signal format and compare them bit by bit.
For different simulation tools, SmootHDL use the same file to generate netlist files through different FPGA logic synthesis compilers, and use the same stimulus file to simulate them to compare their outputs. Formula (11) represents the execution of simulation $\mathcal{S}_{j}$ on netlist $\mathcal{N}$ using stimuli $\sigma$, resulting in output $\mathcal{O}_{j}$. Each simulation tool might use a slightly different engine or algorithm, and $\mathcal{O}_{j}$ captures the output of the $j$-th tool.
@@
After simulations, outputs are compared to assess the consistency of the netlist’s behavior across different simulation platforms based on Formula (12). This formula ensures that all outputs $\mathcal{O}_{i}$ are identical for every pair of simulations, confirming that the netlist behaves consistently irrespective of the simulation tool used. If the output is inconsistent, SmootHDL will report the issue. We will manually check the issues reported by SmootHDL and report to developers.
@@
For comparing original and variant model outputs, SmootHDL prints the signal as a 0/1 electrical signal, and compares the simulation result output after the original Simulink Model is converted to HDL code and synthesized bit by bit with the simulation result output after the variant Simulink Model is converted into HDL code and synthesized. Theoretically the results should not change since the inserted module will not be executed and translated. But if the results are inconsistent, SmootHDL will report the Bugs. We will manually check these issues to find the root causes and submit them to the corresponding software developers.
@@ IV Evaluation
In this section, four experiments are conducted to evaluate the effectiveness of SmootHDL. Specifically, our evaluation aims at answering the following Research Questions (RQs).
RQ1: How is the bug-finding capability of SmootHDL compared to the state-of-the-art methods?
RQ2: Can SmootHDL detect new FPGA logic synthesis compiler bugs?
RQ3: How diverse is the HDL code generated from SmootHDL compared to the state-of-the-art methods?
RQ4: How effective is the BDS of SmootHDL?
In our experiments, RQ1 and RQ2 are used to evaluate the bug-finding capability of SmootHDL compared to the state-of-the-art approaches. RQ3 and RQ4 are employed to evaluate the code generation strategy of the code generation component and the diverse variant generation component.
@@ IV-B Evaluation Setup
SmootHDL has been developed using MATLAB and Python, and both our code and experimental data are publicly accessible on GitHub [10]. The evaluation of SmootHDL was conducted on a computer running the Ubuntu 22.04 operating system, equipped with an Intel Core i9 CPU @ 2.10GHz, and 128GB of memory. Based on the recommendations of existing work, we focus on the size of generated HDL files being in the 700-100 line range [9].
@@ IV-C Answer to RQ1
Approach. To evaluate the effectiveness of SmootHDL, we compare the bug-finding capability of SmootHDL with the state-of-the-art methods Verismith [9] and LegoHDL [7], since finding more bugs within a time period is the main objective of these approaches. In the experiment, we detect bugs on both the recently released version of FPGA logic synthesis compiler include Vivado, yosys, Iverilog and Quartus. We set a single testing period of two weeks for FPGA logic synthesis compiler, that is, every method tests these FPGA synthesis compilers for two weeks.
TABLE I: Bugs detected by Our method, LegoHDL, and Verismith
@@
Our Method 2 4 6
Results. As illustrated in Table I, bugs detected in our experiment are categorized into new bugs (New) and previously identified ones (Known), with the latter being duplicates from the bug repository. Notably, Table I demonstrates that SmootHDL significantly outperforms Verismith and LegoHDL in bug detection capabilities. Over four weeks, SmootHDL identified six bugs in these FPGA logic synthesis compilers, comprising two new bugs and four known ones. In contrast, LegoHDL detected three bugs—two known and one new—while Verismith identified only one known bug. This occurs because SmootHDL is capable of generating HDL test cases that are more complex than those produced by the baseline method. During the translation of the Simulink Model to HDL code by the HDL coder, each subsystem is converted into a separate HDL file with defined reference relationships. SmootHDL enhances the complexity of these relationships by integrating complex code branches into the Simulink Model, thereby challenging the FPGA logic synthesis compiler to employ more sophisticated optimization strategies for the HDL files. Additionally, the Bayesian selector intensively tests the compiler by selecting increasingly complex HDL code scenarios for processing. All these lead to SmootHDL’s bug discovery efficiency being better than the baseline method.
Conclusion. SmootHDL shows better bug-finding capability than LegoHDL and Verismith for testing FPGA logic synthesis compiler. In two week testing period SmootHDL found 6 bugs, which 4 is unknown before.
@@ IV-D Answer to RQ2
Figure 5: FPGA Logic Synthesis Compiler Bug Found via Our Method
Approach. Over a two-month period from May to July 2024, we extensively tested SmootHDL to assess its bug-finding capabilities. Unfortunately, we could not validate the latest version of Quartus due to licensing issues. However, we employed the latest versions of other FPGA logic synthesis compilers for testing. Each identified issue with a logic synthesis compiler was reported to the corresponding software community, and we engaged with developers there to gather feedback.
Results. As depicted in Table II, our approach successfully identified 16 bugs within a three-month timeframe. The official websites of various logic synthesis tools, including Iverilog, Vivado, Yosys, and Quartus, have classified each issue as either Confirmed as New, Confirmed as Known, or Unconfirmed in Doubt (Pending Verification). New indicates bugs previously unrecognized by developers. Known refers to issues already identified as bugs by developers. Pending represents issues that developers believe can be circumvented through specific standardized procedures. To facilitate replication of our findings, we have made the HDL code files that triggered these bugs available on GitHub [10]. Notably, four new bugs were detected that had not been identified by previous methods. Figure5 illustrates one such case revealed by SmootHDL. In this example, our method duplicated a feasible state and inserted an inhibited branch, which after Simulink-to-HDL translation was encapsulated into a new subsystem. The resulting HDL introduced a redundant WideOr6 logic component connected to multiple multiplexers. Although the inhibited guard should prevent this path from being activated, Vivado incorrectly retained it during optimization. As a result, the synthesis tool attempted to propagate signals through this unused logic cone, which involved reset-related nets, eventually leading to a failure in the optimization stage. Conventional methods such as Verismith and LegoHDL are unable to expose this Bugs because they rarely generate designs with such redundant and inhibited branches at the control-flow level. In contrast, SmootHDL’s structural mutations and Bayesian-guided selection deliberately increase both control-flow and timing complexity, making it more effective at revealing hidden synthesis bugs of this kind.
TABLE II: Bugs found by Our Method
@@
Vivado-2 Synthesis failed and crash(HARTGLADDGen:regenerate(bool)) Verified C Vivado
Vivado-3 Vivado synthesis failed (ConstPropagate) Verified C Vivado
@@
Vivado-7 Reference relationship error caused synthesis failed Pending C Vivado
@@
    There are two types of status feedback from the official bug report (i.e., $Verified$ = newly confirmed bug, pendingpending = Pending verification). There are two types of bugs (i.e., Type) in our reported bugs: crash bugs ($C$) and miscompilation bugs ($M$).
Conclusion. SmootHDL has demonstrated effectiveness in detecting bugs in FPGA logic synthesis compilers. Within a two-month period, it identified 16 valid bugs, among which 4 were confirmed as new by the official developers.
@@ IV-E Answer to RQ3
Approach. We employ SmootHDL, LegoHDL, and Verismith to generate HDL code test cases, each consisting of 700–1000 lines and totaling 1000 test cases. We compare their complexity and generation time using three key indicators derived from the Abstract Syntax Tree (AST) and synthesis reports. The first metric, the count of blocks, tallies the blocks in the generated files, where a higher block count reflects greater structural complexity. The second metric, the count of lines or connections, indicates that a higher number of connections increases the likelihood of encountering synthesis issues. The third metric is the critical path delay obtained from synthesis timing analysis, where larger delay values represent deeper logic paths and more complex timing structures that can stress synthesis tools more thoroughly.
Figure 6: Blocks, connections and references generated by Our Method, LegoHDL and Verismith
Results. Numbers of Blocks and Connections. Blocks and connections are vital components of HDL code, crucial for depicting its characteristics. As shown in Figure6, SmootHDL generates a range of [40,49] blocks and [125,226] connections, surpassing the ranges produced by Verismith [9] ([25,37] blocks and [102,117] connections) and LegoHDL ([30,41] blocks and [120,210] connections). This increased complexity arises as SmootHDL continuously selects more intricate HDL codes for testing, based on a Bayesian selector, enhancing its bug detection efficiency over the baseline methods. Delay. Figure6 reports the distribution of critical path delay for LegoHDL, Verismith, and our method. The median delay of LegoHDL is 22.0, while Verismith is lower at 14.5. Our method shows a higher median of 27.6. In terms of range, LegoHDL spans 14–29, Verismith 7–26, and our method 18–35. At the upper end, our method produces a larger proportion of high-delay test cases. Specifically, 66% of cases exceed 25s and 32% exceed 30s, compared with 14% and none for LegoHDL, and 2% and none for Verismith. At the lower end, 75% test cases of Verismith and 27% test cases of LegoHDL fall below 20s, while our method shows only 11% in this range. These observations indicate that, the control-flow mutations and Bayesian-guided selection shift the delay distribution upward, producing HDL test cases with deeper logic paths and therefore higher timing complexity than the two baselines.
Conclusion. Compared with Verismith and LegoHDL, SmootHDL generates HDL test cases with higher structural complexity and richer reference relationships. These results indicate that SmootHDL can provide greater diversity in generated test cases, which is valuable for exercising FPGA logic synthesis compilers.